lf x, y, z are consecutive odd integers and their sum is 27 . What is their product xyz?
A 720 b 693 c 990 d 1287 e 1485

Answer by dabanfield(803)

lf x, y, z are consecutive odd integers and their sum is 27 . What is their product xyz?
A 720 b 693 c 990 d 1287 e 1485
Since the integers are consecutive and odd we know y = x+2 and z = x+4 so we have:
x + (x+2) + (x+4) = 27
3x + 6 = 27
3x = 21
x = 7
y = x+2 = 7+2 = 9
z = x+4 = 7+4 = 11
The product then is 7*9*11 = ?